10                                    20         30 1tum T0090  MKKLQ........IAVGIIRNE...................N.........NEIFITRRAADAHM.A... MLKPDnlpvtfgkNDVEIIAREtlyrgffsldlyrfrhrlf-ngqmshevrREIFERGHAAVLLPf-dpv 40        50        60 1tum T0090  .....................NKLEFPGGKIEMGETPEQAVVRELQEEVGITPQH....FSLF....... rdevvlieqiriaaydtsetpWLLEMVAGMIEEGESVEDVARREAIEEAGLIVKRtkpvLSFLaspggts 70                         80        90       100       110     120 1tum T0090  .EKLE.................YEFPDRHITLWFWLVERWEGEPWGKEGQPGEWMSLVGLNADDFPPANE e----rssimvgevdattasgiHGLADENEDIRVHVVSREQAYQWVEEGKIDNAASVIALQWLQLHH--- 1tum T0090  PVIAKLKRL -QALKNEWA

Figure 1: Comparison of candidate test.a2m.gz (lines) to reference 1tum-T0090-fssp-global-fw0.5.a2m.gz (letters). Reference alignment has 119 matching columns, Candidate alignment has 124 matching columns. 24 were used (for template shift+target shift) to compute shift score. The shift score is 0.000000, with a min shift of 114, max shift of 0, and mean of 0.000000. Optimal subalignment has 114 matching columns, with 119 paired matches. 0 were used (for template shift+target shift) to compute shift score. Optimum shift score is 0.935969, with a min shift of 233, max shift of 0, and mean of 0.407222. Optimum achieved by removing the 15 columns indicated with dotted lines from the candidate alignment.